OpenStudy (anonymous):

the length of a rectangular photograph is 7in more than width. if the area is 78in, what are the dimensions of the photograph?

OpenStudy (anonymous):

(x+7)(x)=78

OpenStudy (anonymous):

x^2+7x=78

OpenStudy (watchmath):

L=length W=width L=W+7 LW=78 (W+7)W=78 W^2+7W-78=0 (W+13)(W-6)=0 W=-17(can't be negative) or W=6 (this is the one) So L=6+7=13 Dimension = L times W = 13 by 6

OpenStudy (anonymous):

thank you =]